a teacher gives four students their written papers randomly
what is the probability that :
1- each one gets his own paper?
2- the first one gets his own paper?
3- 2 of them get their own papers?
4- no one gets his own paper ?

Answer by edjones(8007)

1) 1/4!=1/24
2) 1/4
3) 2 of them get their own papers (and 2 do not): 6/24=1/4
4) 9/24=3/8
For 3 and 4 I enumerated the possible permutations in order to find the answers.
.
Ed